$SPAD/src/input poly.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool poly.output
)set message test on
)set message auto off
)clear all
--S 1 of 54
a := rootOf(a**4+1,a)
--R
--R
--R   (1)  a
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 54
definingPolynomial a
--R
--R
--R         4
--R   (2)  a  + 1
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 54
b := rootOf(b**2-a-1,b)
--R
--R
--R   (3)  b
--R                                                    Type: Expression(Integer)
--E 3
--S 4 of 54
a + b
--R
--R
--R   (4)  b + a
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 54
% ** 5

--R
--R
--R            3      2                 3      2
--R   (5)  (10a  + 11a  + 2a - 4)b + 15a  + 10a  + 4a - 10
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 54
rootOf(c**2+c+1,c)
--R
--R
--R   (6)  c
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 54
zeroOf(d**2+d+1,d)
--R
--R
--R         +---+
--R        \|- 3  - 1
--R   (7)  ----------
--R             2
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 54
rootOf(e**5-2,e)
--R
--R
--R   (8)  e
--R                                                    Type: Expression(Integer)
--E 8
--S 9 of 54
zeroOf(f**5-2,f)
--R
--R
--R        5+-+
--R   (9)  \|2
--R                                                    Type: Expression(Integer)
--E 9
)clear all
--S 10 of 54
p := 3*x**8 + 2*x**7 + 6*x**2 + 7*x + 2
--R
--R
--R          8     7     2

--R   (1)  3x  + 2x  + 6x  + 7x + 2
--R                                                    Type: Polynomial(Integer)
--E 10
--S 11 of 54
q := 2*x**13 + 9*x**7 + 2*x**6 + 10*x + 5
--R
--R
--R          13     7     6
--R   (2)  2x   + 9x  + 2x  + 10x + 5
--R                                                    Type: Polynomial(Integer)
--E 11
--S 12 of 54
gcd(p,q)
--R
--R
--R         7
--R   (3)  x  + 2x + 1
--R                                                    Type: Polynomial(Integer)
--E 12
--S 13 of 54
resultant(p,q,x)
--R
--R
--R   (4)  0
--R                                                    Type: Polynomial(Integer)
--E 13
)clear all
--S 14 of 54
p := x**2 + y**2
--R
--R
--R         2    2
--R   (1)  y  + x
--R                                                    Type: Polynomial(Integer)
--E 14
--S 15 of 54
eval(p,x=5)
--R
--R
--R         2
--R   (2)  y  + 25
--R                                                    Type: Polynomial(Integer)
--E 15

--S 16 of 54
eval(p,[x = a + b,y = c + d])
--R
--R
--R         2           2    2           2
--R   (3)  d  + 2c d + c  + b  + 2a b + a
--R                                                    Type: Polynomial(Integer)
--E 16
--S 17 of 54
q := x**3 + 5*x - y**4
--R
--R
--R           4    3
--R   (4)  - y  + x  + 5x
--R                                                    Type: Polynomial(Integer)
--E 17
--S 18 of 54
eval(q,[x=y,y=x])
--R
--R
--R         3         4
--R   (5)  y  + 5y - x
--R                                                    Type: Polynomial(Integer)
--E 18
--S 19 of 54
px := eval(p, y = sin(2.0))
--R
--R
--R         2
--R   (6)  x  + 0.8268218104 3180595732
--R                                                      Type: Polynomial(Float)
--E 19
--S 20 of 54
eval(px, x = cos(2.0))
--R
--R
--R   (7)  1.0
--R                                                      Type: Polynomial(Float)
--E 20
)clear all
--S 21 of 54
factor(x**3 - 3*x + 2)
--R
--R

--R               2
--R   (1)  (x - 1) (x + 2)
--R                                          Type: Factored(Polynomial(Integer))
--E 21
--S 22 of 54
factor(x**2/4 + x*y + y**2)
--R
--R
--R             1   2
--R   (2)  (y + - x)
--R             2
--R                                Type: Factored(Polynomial(Fraction(Integer)))
--E 22
--S 23 of 54
p := x**3 + x*y + 2*x**2*y**2 + 2*y**3 + 3*x**2*z + 6*x*y**2*z
--R
--R
--R             2     2       3     2 2          3
--R   (3)  (6x y  + 3x )z + 2y  + 2x y  + x y + x
--R                                                    Type: Polynomial(Integer)
--E 23
--S 24 of 54
factors := factor p
--R
--R
--R           2                  2
--R   (4)  (2y  + x)(3x z + y + x )
--R                                          Type: Factored(Polynomial(Integer))
--E 24
--S 25 of 54
nthFactor(factors,1)
--R
--R
--R          2
--R   (5)  2y  + x
--R                                                    Type: Polynomial(Integer)
--E 25
--S 26 of 54
nthFactor(factors,2)
--R
--R
--R                    2
--R   (6)  3x z + y + x
--R                                                    Type: Polynomial(Integer)
--E 26

)clear all
--S 27 of 54
p := a*x**2 + b*x*y + c*y**2
--R
--R
--R           2              2
--R   (1)  c y  + b x y + a x
--R                                                    Type: Polynomial(Integer)
--E 27
--S 28 of 54
q := 13*x**2 + 3*z
--R
--R
--R                2
--R   (2)  3z + 13x
--R                                                    Type: Polynomial(Integer)
--E 28
--S 29 of 54
p + q
--R
--R
--R                2                    2
--R   (3)  3z + c y  + b x y + (a + 13)x
--R                                                    Type: Polynomial(Integer)
--E 29
--S 30 of 54
p - 3*q
--R
--R
--R                  2                    2
--R   (4)  - 9z + c y  + b x y + (a - 39)x
--R                                                    Type: Polynomial(Integer)
--E 30
--S 31 of 54
p**2 + p*q
--R
--R
--R   (5)
--R          2                2      2 4           3                  2  2 2
--R     (3c y  + 3b x y + 3a x )z + c y  + 2b c x y  + ((2a + 13)c + b )x y
--R   +
--R                 3      2        4
--R     (2a + 13)b x y + (a  + 13a)x
--R                                                    Type: Polynomial(Integer)

--E 31
--S 32 of 54
r := (p + q)**2
--R
--R
--R   (6)
--R       2        2                      2      2 4           3
--R     9z  + (6c y  + 6b x y + (6a + 78)x )z + c y  + 2b c x y
--R   +
--R                    2  2 2               3      2              4
--R     ((2a + 26)c + b )x y  + (2a + 26)b x y + (a  + 26a + 169)x
--R                                                    Type: Polynomial(Integer)
--E 32
--S 33 of 54
setVariableOrder [a,b,c,x,y,z]
--R
--R                                                                   Type: Void
--E 33
--S 34 of 54
p
--R
--R
--R         2             2
--R   (8)  x a + y x b + y c
--R                                                    Type: Polynomial(Integer)
--E 34
--S 35 of 54
q
--R
--R
--R           2
--R   (9)  13x  + 3z
--R                                                    Type: Polynomial(Integer)
--E 35
--S 36 of 54
r
--R
--R
--R   (10)
--R      4 2        3      2 2       4       2      2 2 2
--R     x a  + (2y x b + 2y x c + 26x  + 6z x )a + y x b
--R   +
--R      3           3               4 2       2 2       2         4        2     2
--R   (2y x c + 26y x  + 6z y x)b + y c  + (26y x  + 6z y )c + 169x  + 78z x  + 9z
--R                                                    Type: Polynomial(Integer)

--E 36
--S 37 of 54
resetVariableOrder()
--R
--R                                                                   Type: Void
--E 37
--S 38 of 54
p
--R
--R
--R            2              2
--R   (12)  c y  + b x y + a x
--R                                                    Type: Polynomial(Integer)
--E 38
--S 39 of 54
coefficient(q,x,2)
--R
--R
--R   (13)  13
--R                                                    Type: Polynomial(Integer)
--E 39
--S 40 of 54
coefficient(r,x,3)
--R
--R
--R   (14)  (2a + 26)b y
--R                                                    Type: Polynomial(Integer)
--E 40
--S 41 of 54
c := coefficient(r,z,1)
--R
--R
--R             2                      2
--R   (15)  6c y  + 6b x y + (6a + 78)x
--R                                                    Type: Polynomial(Integer)
--E 41
--S 42 of 54
coefficient(c,x,2)
--R
--R
--R   (16)  6a + 78
--R                                                    Type: Polynomial(Integer)
--E 42

--S 43 of 54
coefficient(q**2, [x,z], [2,1])
--R
--R
--R   (17)  78
--R                                                    Type: Polynomial(Integer)
--E 43
--S 44 of 54
coefficient(r, [x,y], [2,2])
--R
--R
--R                       2
--R   (18)  (2a + 26)c + b
--R                                                    Type: Polynomial(Integer)
--E 44

)clear all
--S 45 of 54
l := rootsOf(x**4+1,x)
--R
--R
--R   (1)  [%x0,%x0 %x1,- %x0,- %x0 %x1]
--R                                              Type: List(Expression(Integer))
--E 45
--S 46 of 54
x0**5
--R
--R
--R          5
--R   (2)  x0
--R                                                    Type: Polynomial(Integer)
--E 46
--S 47 of 54
definingPolynomial x0
--R
--R
--R   (3)  - x0 + %%var
--R                                                    Type: Expression(Integer)
--E 47
--S 48 of 54
definingPolynomial x1
--R
--R
--R   (4)  - x1 + %%var

--R                                                    Type: Expression(Integer)
--E 48
--S 49 of 54
definingPolynomial x2
--R
--R
--R   (5)  - x2 + %%var
--R                                                    Type: Expression(Integer)
--E 49
--S 50 of 54
x3 := last l
--R
--R
--R   (6)  - %x0 %x1
--R                                                    Type: Expression(Integer)
--E 50
--S 51 of 54
x0 + x1 + x2 + x3
--R
--R
--R   (7)  - %x0 %x1 + x2 + x1 + x0
--R                                                    Type: Expression(Integer)
--E 51
--S 52 of 54
x0 * x1 * x2 * x3
--R
--R
--R   (8)  - x0 x1 x2 %x0 %x1
--R                                                    Type: Expression(Integer)
--E 52
--S 53 of 54
zerosOf(y**4+1,y)
--R
--R
--R          +---+      +---+        +---+        +---+
--R         \|- 1  + 1 \|- 1  - 1 - \|- 1  - 1 - \|- 1  + 1
--R   (9)  [----------,----------,------------,------------]
--R             +-+        +-+         +-+          +-+
--R            \|2        \|2         \|2          \|2
--R                                              Type: List(Expression(Integer))
--E 53
--S 54 of 54
definingPolynomial y1
--R

--R
--R   (10)  - y1 + %%var
--R                                                    Type: Expression(Integer)
--E 54
)spool
)lisp (bye)

———-

References^{[1] nothing}

14